40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Tue Dec 4 19:34:22 2018, awade, DailyProgress, WOPO, Diabolo 532 nm power drop and retuning 18-12-04_DiaboloInside_annotate.jpegplot20181204_DiaboloSHGRaligned1064nm.pdfplot20181204_DiaboloSHGRErrSig.pdf20181204_RetuneDiaboloSHGAlignment.zip
    Reply  Thu Dec 6 20:04:50 2018, awade, DailyProgress, WOPO, Diabolo 532 nm power drop and retuning 
       Reply  Fri Dec 7 20:31:06 2018, awade, DailyProgress, WOPO, Diabolo: cleaning dust out of SHG unit, self heating issues 18-12-07_20-06-46_4774.jpg18-12-07_20-06-58_4776.jpg18-12-07_20-07-44_4780.jpgIMG_4785.JPGIMG_4766.MOV
          Reply  Thu Dec 13 14:01:13 2018, awade, DailyProgress, WOPO, Diabolo: contacted coherent 
             Reply  Thu Dec 13 16:24:06 2018, awade, DailyProgress, WOPO, Diabolo: baseline numbers for coherent plot20181218_PhaseMatchingCurveFastScanDiaboloSHG.pdf20181213_SHGScanWithLNTempTuning.zip
                Reply  Wed Dec 19 16:48:18 2018, awade, DailyProgress, WOPO, Diabolo: baseline numbers for coherent 
                   Reply  Thu Dec 20 19:20:50 2018, awade, DailyProgress, WOPO, Diabolo: trialing SHG oven set temperature points to see if there is a maxima when locked plot20181220_SHGLockedOutputVsOvenSetpointTemperature.pdf20181220_SHGLockedPowerVsSetPointTemp.zip
                      Reply  Fri Dec 21 15:33:26 2018, Koji, DailyProgress, WOPO, Diabolo: trialing SHG oven set temperature points to see if there is a maxima when locked 
                         Reply  Fri Dec 21 15:51:27 2018, awade, DailyProgress, WOPO, Diabolo: trialing SHG oven set temperature points to see if there is a maxima when locked 
                      Reply  Mon Dec 24 10:08:49 2018, awade, DailyProgress, WOPO, Diabolo: back to power with deliberate pre-misalignment 18-12-21_22-04-30_4812.jpg18-12-21_22-04-54_4813.jpg18-12-21_22-07-55_4814.jpg
                         Reply  Mon Dec 24 12:07:18 2018, Koji, DailyProgress, WOPO, Diabolo: back to power with deliberate pre-misalignment 
Message ID: 2273     Entry time: Thu Dec 6 20:04:50 2018     In reply to: 2268     Reply to this: 2277
Author: awade 
Type: DailyProgress 
Category: WOPO 
Subject: Diabolo 532 nm power drop and retuning 

I've continued to have problems getting the Diabolo laser SHG to lock. 

I get nice clean fundamental resonance peaks when fast scanning the SHG cavity length and optimizing the 1064 nm alignment into the SHG . However, when I switch to engage I see an initial power jump before the power dips and then settles to what looks like a higher order mode.  It seems like the heating of the NL crystal when locked is misaligning the eigen-mode of the cavity so that the cold scan of the cavity is optimal until locking and the higher circulating power heats things up.  

I tried a different approach, slow scanning close about the fundamental (to keep things warm) and attempted to maximize the fundamental peak.   This was done by narrowing the scan range and then tuning the laser frequency (using NPRO crystal temp) to be close to resonance.  This improved the power output by a few tens of percent to ~27 mW, but is still well short of the previous performance.  This current power output is unworkable because we need at least 60 mW to overcome losses coupling into the WOPO and also because it is not stable in power and drifts around a bit and sometimes unlocks (probably because it is on the edge of some thermal state of the system).

I also tried tuning 1064 nm alignment into the cavity with the SHG locked.  This is tricky because that for each adjustment it takes a little time for the thermal state of the crystal to stabilize and settle. It improved power output maybe 10% but wasn't clearly leading to good operating point quickly.

Fixing this

So it seems the unit was previously in a combination of states that worked once the SHG was locked.  Its been hard to reproduce that operating point. Its possible that since 2004, when the unit was manufactured, that there has been some degradation to the crystal.  Also, maybe the reason it has failed to perform to spec might be that the SHG cavity itself is slightly mialigned.  I'm reluctant to open up the SHG unit as it is hermetically sealed.  I assume they flood it with N2 to keep everything in a dry no oxygen controled environment but at this stage maybe it doesn't matter.  Tweaking the input coupler on the SHG will be super sensitive so would be a delecate task, but I'm at a loss as to how to find an optimal operating point to return to normal operation.

 

ELOG V3.1.3-